If the heat of combustion for a specific compound is - 1200.0 kJ/mol and its molar mass is 29.71 g/mol, how many grams of this compound must you burn to release 891.80 kJ of heat?

 mass: _______ g
Answer #1

1200 KJ heat is released by 1 mole compound

So, 891.80 KJ heat will be released by = 891.80/1200 = 0.7432 moles of compound

Mass of compound = moles * molar mass = 0.7432 * 29.71 = 22.08 grams